OpenStudy (anonymous):

Dear All, What is the simplest method to solve the simultaneous equation below:- Y=-3x+2 Y=2x^2-7x-4 Thanks.

OpenStudy (lgbasallote):

equate both sides \[-3x + 2 = -2x^2 - 7x - 4\] since y = y

OpenStudy (anonymous):

substitution xD

OpenStudy (anonymous):

Thankyou, Best regards to you all

Directrix (directrix):

Simplified from above: -2x^2 +4x +6 =0